OpenStudy (marg):

Highway safety engineers build soft barriers so that cars hitting them will slow down at a safe rate. A person wearing a safety belt can withstand an acceleration of 300 m/s 2 . How thick should barriers be to safely stop a car that hits the barriers at 106 km/h? Answer in units of m

OpenStudy (anonymous):

I assume -3.0 102 m/s2 means -3.0*10^2 m/s^2 which is -300 m/s^2. First, we'll convert 109 km/hr to m/s just to get everything at the same unit: (109 km / 1 hr)(1000 m / 1 km)(1 hour / 3600 sec) = 30.28 m/s Now, we can get the answer to this problem with one step using the equation: v^2 = vo^2 + 2ad, where v = final velocity vo = initial velocity a = acceleration d = displacement Now, we know the car's initial velocity is 30.28 m/s, we know its final velocity is 0 m/s, and we know its acceleration is -300 m/s^2. So, we can plug all of this in and solve for displacement: v^2 = vo^2 + 2ad 0^2 = (30.28)^2 + 2(-300)d d = 1.53 m = answer

OpenStudy (marg):

thank you:)

OpenStudy (anonymous):

your welcome